When parking a car in a downtown parking lot, drivers pay according to the number of hours or fraction thereof. The probability distribution of the number of hours cars are parked has been estimated as follows: 

X12345678
P(X)0.2130.1170.120.0850.0620.0280.0230.352

A. Mean = _______ 

B. Standard Deviation = _______ 

The cost of parking is 3.25 dollars per hour. Calculate the mean and standard deviation of the amount of revenue each car generates 

A. Mean = _______ 

B. Standard Deviation = _______ 
Answer #1
X P(X) Mean= X.P(X) (X-Mx)^2 Variance= P(X).(X-Mx)^2
1 0.213 0.213 12.974404 2.763548052
2 0.117 0.234 6.770404 0.792137268
3 0.12 0.36 2.566404 0.30796848
4 0.085 0.34 0.362404 0.03080434
5 0.062 0.31 0.158404 0.009821048
6 0.028 0.168 1.954404 0.054723312
7 0.023 0.161 5.750404 0.132259292
8 0.352 2.816 11.546404 4.064334208
Sum of X.P(X), Mx 4.602
Sum of P(X).(X-Mx)^2, Vx 8.155596
Standard Deviation 2.855800413

a) Mean, Mx= 4.602

b) SD= sqrt(Vx)= 2.855

3.25 $ per hour

c) Mean revenue= 4.602*3.25= 14.956$

d) SD= 2.855*3.25= 9.278 $

Hope this helps and Please don't forget to rate the answer :)
Answer #2

How did you calculate the standard deviation?

answered by: Jack Freeman